find the domain
y=sqrt 4-x^2

Answer by stanbon(48535)

find the domain
y=sqrt 4-x^2
---------------
The radicand cannot be negative:
4-x^2>=0
x^2-4<=0
(x-2)(x+2)<=0
----------------------
Mark x=-2 and x=+2 on a number line.
Check test values in each of the resulting x-intervals to see where the
solution intervals lie.
Test:
x=-3 ; 9-4<=0 ; false
x=0: 0-4<=0; true
x=3; 9-4<=0; false
-----------------------
Solution interval: (-2,2)
============================
Cheers,
Stan H.